OpenStudy (anonymous):

Can someone explain how to do this? |x + 3| < 5

OpenStudy (darkprince14):

x+3< 5 x+3< -5 find the value of x look for its domain..

OpenStudy (anonymous):

To solve for x you need to get the x by itself. So on the first equation you need to subtract 3 from each side. Which brings you to x<5-3, then solve the 5-3. So you get x<2. The second one is the same.

OpenStudy (amorfide):

that isn't a very good explanation there are two methods a graphical method to do the graphical method draw the graph y=x+3 then draw the graph y=5 then draw the graph y=|x+3| the |x+3| means that when it is about to cross the x axis, it will bounce off and go through positive version of negative points for example x+3 if i substitute x=-4 -4+3=-1 instead of -1, it will go through the point 1 so you get a graph looking like this  so you want to find where the to points intersect you know it intersects twice

OpenStudy (amorfide):

so looking at the graph you know that the line y=x+3 crosses with the line y=5 so you have an equation x+3<5 x<2 then you know that the line y=|x+3| and y=5 cross but |x+3| just means that any negative value, becomes positive, so instead of writing |x+3| you can see that this is x+3 but going in the opposite direction, so this is actually a reflection in the y axis so we write -(x+3) so this gives us another equation -(x+3)<5 -x-3<5 -x<8 x>-8